A farmer has a rectangular garden plot surrounded by 220 feet of fence. Find the dimensions of the garden if its area is 2400 square feet.
Answer by checkley77(12844)

2L+2W=220
LW=2,400 OR L=2,400/W
2(2,400/W)+2W=220
4,800/W+2W=220
(4,800+2W*W)/W=220
(2W^2+4,800)/W=220
2W^2+4,800=220*W
2W^2+4,800=220W
2W^2-220W+4,800=0
2(W^2-110W+2,400)=0
2(W-880)(W-30)=0
W-80=0
W=80 L=2,400/80=30
W-30=0
W=30 L=2,400/30=80
PROOF:
2*30+2*80=220
60+160=220
220=220